what is an ordered triple that would be a solution to the quation
2x-4y+3z=10

Answer by stanbon(75887)

what is an ordered triple that would be a solution to the quation
2x-4y+3z=10
---
Let y = 5 and let z = 2, then solve for "x":
2x -4(5) + 3(2) = 10
2x -14 = 10
2x = 24
x = 12
----
Your triple: (12,5,2)
============================
Cheers,
Stan H.